### v1

@@ 5 Methods
Modelling approach Our modelling approach (Fig. 16) soft links the SBSP model and the PyPSA-Eur [42] model to assess the cost-competitiveness of SBSP for European power system decarbonization in 2020 and 2050. Initially, we model the power generation of various SBSP designs using the methodology outlined in Section 2, incorporating generation profiles and cost data for RD1 and RD2. Next, we run the PyPSA-Eur model to optimize power generation scenarios for 2020 and 2050, performing time-step optimizations in 3-hour intervals. This includes validating the 2020 network as a reference for further sensitivity analyses. Finally, different SBSP designs are integrated into the PyPSA-Eur model as generators, and re-optimized to assess the impact of SBSP on both the current and future European electrical systems, as well as its technical feasibility. Across various scenarios, we analyze system generation costs, total energy generation by different sources, the relationship between major energy sources and SBSP costs, greenhouse gas emissions, and the countries benefiting from SBSP deployment.
@@
Electricity Demand The hourly electricity demand profiles for each country of the network in 2020 are taken from the ENTSO-E website [45]. For the electricity demand projections in 2050 across various countries, we utilize the predictions from the ENTSO-E & ENTSOG TYNDP 2024 Scenarios [46], specifically drawing on the modeling results for electricity, hydrogen, hybrid heat pumps, and synthetic fuels.
@@
Additionally, in our model, the storage module includes batteries, existing pumped-hydro storage (PHS), hydroelectric dams, hydrogen storage, and synthetic energy carriers such as methane.
@@
Figure 17(a) demonstrates that our model aligns closely with existing databases, capturing 993 GW out of the 1048 GW of installed capacity reported by ENTSO-E [56, 57]. Discrepancies in coal, natural gas, oil, and biomass capacities can largely be attributed to the lag in data updates, which omits recently commissioned plants and results in an underestimation of these capacities in our model. Additionally, using ERA5 data to model solar potential introduces a known bias, as ERA5-derived solar capacity tends to be overestimated [58], leading to an inflated solar capacity representation. However, given the model’s inclusion of 33 countries, validation from a technology-only perspective is insufficient to fully assess its accuracy. Figure 18 therefore provides additional validation at the country level, showing that most countries align with reported data to a high degree of accuracy (2%–15% error). Apart from installed capacity, Figure 17(b) shows that most generation types have comparable power outputs (see Table 4 for numerical data). Beyond minor deviations due to capacity differences, the main variation appears in the generation mix, particularly between coal and natural gas. Higher input marginal prices for natural gas relative to coal result in lower natural gas generation. Looking ahead to 2050 scenarios, as carbon emissions are set to zero, fossil-based sources such as coal, natural gas, and oil will be phased out, rendering this discrepancy non-influential for future analyses.
@@
Soft-linked SBSP-PyPSA-Eur model Among the two integrated models, the SBSP model emphasizes simulated parameters for two representative designs, RD1 and RD2, focusing on annual generation profiles and associated costs projected for 2020 and 2050. The PyPSA-Eur model, on the other hand, serves as an electricity system optimization tool that aims to minimize total annualized system costs across various scenarios.
Using RD1 as an example, we incorporate SBSP as a new generator type into the updated 2020 and 2050 networks. Given that a single SBSP satellite can deliver energy to multiple regions simultaneously [18], we assume each node has a dedicated ground station to receive this energy. Thus, an SBSP generator is added at each node, simulating the simultaneous energy reception across nodes. We set the capacity parameter p_nom as extendable, allowing it to adjust according to demand. Additionally, p_max_pu is updated to align with the SBSP generation profile data, normalized for consistency.
At the network level in 2020 and 2050, all generator capacities are set as extendable, allowing generation to be governed by cost-efficiency. However, to enhance model realism, we impose specific constraints on different generator types. Due to geographical and environmental limitations, run-of-river and reservoir generators have their maximum capacity p_nom_max restricted to current installed capacity p_nom, reflecting the limited expansion potential [59, 60]. Biomass and nuclear generators are assigned a minimum capacity p_nom_min equivalent to their current capacity p_nom, acknowledging the high construction and operational costs of nuclear plants and the sustainable, carbon-neutral attributes of biomass energy [61, 62]. Generators powered by coal, oil, and natural gas, projected to be phased out, have their p_nom set to 0.
@@ Acknowledgements
We acknowledge technical discussions with Gabriel Wittenberg. W.H. would like to acknowledge the support of the Royal Academy of Engineering (RAEng) Engineering for Development Research Fellowship [grant number RF/201819/18/89] and UKRI-EPSRC [grant number EP/W027372/1].

### v2

@@ 5 Methods
Modelling approach Our modelling approach (Fig. 8) soft links the SBSP model and the PyPSA-Eur [56] model to assess the cost-competitiveness of SBSP for European power system decarbonization in 2020 and 2050. Initially, we model the power generation of various SBSP designs using the methodology outlined in Section 2, incorporating generation profiles and cost data for heliostat design and planar design. Next, we run the PyPSA-Eur model to optimize power generation scenarios for 2020 and 2050, performing time-step optimizations in 3-hour intervals. This includes validating the 2020 network as a reference for further sensitivity analyses. Finally, different SBSP designs are integrated into the PyPSA-Eur model as generators, and re-optimized to assess the impact of SBSP on both the current and future European electrical systems, as well as its technical feasibility. Across various scenarios, we analyze system generation costs, total energy generation by different sources, the relationship between major energy sources and SBSP costs, greenhouse gas emissions, and the countries benefiting from SBSP deployment.
@@
Electricity Demand In this study, electricity demand refers exclusively to electricity consumption without including the electrification of transport, heating, or other emerging sectors, both in the 2020 and 2050 scenarios. The hourly electricity demand profiles for each country in 2020 are sourced from the ENTSO-E website [59]. For 2050, we adopt the “Native Demand” projections from the ENTSO-E & ENTSOG TYNDP 2024 Scenarios [22] (Global Ambition, climate year 2009), which represent electricity demand from centralized market, residential, and tertiary sectors, while explicitly excluding additional demands associated with battery charging, electric vehicles (EVs), heat pumps, electrolyzers, and other sector-coupled loads.
@@
Additionally, in our model, the storage module includes batteries, existing pumped-hydro storage (PHS), hydroelectric dams, hydrogen storage, and synthetic energy carriers such as methane. It is worth noting that although hydrogen has potential applications across multiple sectors, in this study, hydrogen storage is modeled solely for electricity balancing purposes, as our analysis focuses exclusively on the electricity sector.
@@
Figure 9(a) demonstrates that our model aligns closely with existing databases, capturing 993 GW out of the 1048 GW of installed capacity reported by ENTSO-E [69, 70]. Discrepancies in coal, natural gas, oil, and biomass capacities can largely be attributed to the lag in data updates, which omits recently commissioned plants and results in an underestimation of these capacities in our model. Additionally, using ERA5 data to model solar potential introduces a known bias, as ERA5-derived solar capacity tends to be overestimated [71], leading to an inflated solar capacity representation. However, given the model’s inclusion of 33 countries, validation from a technology-only perspective is insufficient to fully assess its accuracy. Figure 10 therefore provides additional validation at the country level, showing that most countries align with reported data to a high degree of accuracy (2%–15% error). Apart from installed capacity, Figure 9(b) shows that most generation types have comparable power outputs (see Table 3 for numerical data). Beyond minor deviations due to capacity differences, the main variation appears in the generation mix, particularly between coal and natural gas. Higher input marginal prices of natural gas compared to coal, combined with the absence of a CO2 pricing mechanism, result in lower natural gas generation. Looking ahead to 2050 scenarios, as carbon emissions are set to zero, fossil-based sources such as coal, natural gas, and oil will be phased out, rendering this discrepancy non-influential for future analyses.
@@
Soft-linked SBSP-PyPSA-Eur model Among the two integrated models, the SBSP model emphasizes simulated parameters for two representative designs, heliostat design and planar design, focusing on annual generation profiles and associated costs projected for 2020 and 2050. The PyPSA-Eur model, on the other hand, serves as an electricity system optimization tool that aims to minimize total annualized system costs across various scenarios.
Using heliostat design as an example, we incorporate SBSP as a new generator type into the updated 2020 and 2050 networks. Given that a single SBSP satellite can deliver energy to multiple regions simultaneously [23], we assume each node has a dedicated ground station to receive this energy. Thus, an SBSP generator is added at each node, simulating the simultaneous energy reception across nodes. We set the capacity parameter p_nom as extendable, allowing it to adjust according to demand. Additionally, p_max_pu is updated to align with the SBSP generation profile data, normalized for consistency. This modeling approach reflects the treatment of SBSP in a similar way to other variable renewable sources such as solar and wind, where generation availability varies over time. Within these time-dependent constraints of availability, SBSP is considered fully dispatchable on an hourly basis, allowing its output to respond flexibly to the needs of the system.
At the network level in 2020 and 2050, all generator capacities are set as extendable, allowing generation to be governed by cost-efficiency. However, to enhance model realism, we impose specific constraints on different generator types. Due to geographical and environmental limitations, run-of-river and reservoir generators have their maximum capacity p_nom_max restricted to current installed capacity p_nom, reflecting the limited expansion potential [72, 73]. Biomass and nuclear generators are assigned a minimum capacity p_nom_min equivalent to their current capacity p_nom, acknowledging the high construction and operational costs of nuclear plants and the sustainable, carbon-neutral attributes of biomass energy [74, 75, 76]. Generators powered by coal, oil, and natural gas, projected to be phased out, have their p_nom set to 0.
@@ Acknowledgements
We acknowledge technical discussions with Gabriel Wittenberg. W.H. would like to acknowledge the support of the Royal Academy of Engineering (RAEng) Engineering for Development Research Fellowship [grant number RF
201819
18
89] and UKRI-EPSRC [grant number EP/W027372/1].